# Detecting Functional Bugs in Smart Contracts through LLM-Powered and Bug-Oriented Composite Analysis

### v1

@@ Abstract.
To address these challenges, we design and implement PromFuzz, an automated and scalable system to detect functional bugs, in smart contracts. In PromFuzz, we first propose a novel Large Language Model (LLM)-driven analysis framework, which leverages a dual-agent prompt engineering strategy to pinpoint potentially vulnerable functions for further scrutiny. We then implement a dual-stage coupling approach, which focuses on generating invariant checkers that leverage logic information extracted from potentially vulnerable functions. Finally, we design a bug-oriented fuzzing engine, which maps the logical information from the high-level business model to the low-level smart contract implementations, and performs the bug-oriented fuzzing on targeted functions. We evaluate PromFuzz from $4$ perspectives on $6$ ground-truth datasets and compare it with multiple state-of-the-art methods. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, marking at least a $50\%$ improvement in both metrics over state-of-the-art methods. Moreover, we perform an in-depth analysis on $6$ real-world DeFi projects and detect $30$ zero-day bugs. Our further case studies, the risky first deposit bug and the AMM price oracle manipulation bug on real-world DeFi projects, demonstrate the serious risks of the exploitable functional bugs in smart contracts. Up to now, $24$ zero-day bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.
 † Binbin Zhao and Xingshuang Lin are the co-first authors. * Shouling Ji is the corresponding author.
1. Introduction
@@
Blockchains have increasingly become a crucial component of the contemporary economic landscape. As of the writing of this paper, the total market capitalization of global cryptocurrencies has reached \$$\$3.5$ trillion (for, 2024). Smart contracts, which are fundamental pillars of the blockchain, facilitate the development of decentralized financial (DeFi) and various business transactions. Nevertheless, smart contracts are prone to exploitable bugs that often result in substantial financial losses. According to a report by CertiK, a leading Web3 security firm, over \$$\$1.8$ billion was lost due to $751$ security incidents in 2023 (cer, 2023).
@@
Currently, there are many works have focused on detecting exploitable bugs in smart contracts, which can be categorized into four main types based on their methodologies: static analysis (Feist et al., 2019; Wang et al., 2019b), fuzzing (Choi et al., 2021; Shou et al., 2023), verification (So et al., 2020; Tan et al., 2022), and symbolic execution (myt, 2024; So et al., 2021). Despite these efforts, a recent study (Zhang et al., 2023) reveals that over 80% of exploitable bugs remain undetected by current tools, which are primarily functional bugs. This limitation mainly stems from the tools’ reliance on simple or hard-coded oracles to identify exploitable bugs. Nevertheless, detecting functional bugs involves comprehending the high-level business logic before scrutinizing the low-level implementations in smart contracts. Therefore, there is a pressing need for a practical system that can automatically detect functional bugs in smart contracts.
1.1. Challenges
@@ 1.3. Contributions
$\bullet$ To the best of our knowledge, we build the first ground-truth dataset for verified functional bugs in smart contracts. This dataset includes 23 bugs from various DeFi projects. To facilitate future blockchain security research, we open-source both PromFuzz and the dataset at https://anonymous.4open.science/r/promfuzz-550D.
$\bullet$ We have implemented and evaluated PromFuzz across $4$ dimensions using $6$ ground-truth datasets. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs. PromFuzz demonstrates its superiority in at least a $50\%$ improvement in both metrics over state-of-the-art methods.
$\bullet$ Our extensive analysis of $6$ real-world DeFi projects demonstrates PromFuzz’s great performance in detecting real-world functional bugs. Up to now, PromFuzz has identified $30$ zero-day bugs in $6$ real-world DeFi projects, with $24$ of these bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling \$$\$18.2$ billion from potential monetary losses.
2. Background
@@
2.1. Functional Bugs in Smart Contract
In this paper, we primarily focus on four major categories of functional bugs, including a total of ten subcategories of functional bugs, as shown in Table 1. These classifications cover the majority of functional bugs and are derived from prior research (Sun et al., 2024; Zhang et al., 2023).
@@ 2.1. Functional Bugs in Smart Contract
Table 1. Categories of functional bugs.
Bug Category Bug Subcategory
Price Oracle Manipulation AMM Price Oracle Manipulation
Non-AMM Price Oracle Manipulation
Unauthorized Behavior Approval Not Clear
Unauthorized Transfer
Insecure Calculating Logic Wrong Checkpoint Order
Wrong Interest Rate Order
Risky First Deposit
Incorrect Control Mechanism Improper Handling of the Deposit Fee
Wrong Implementation of Amount Lock
Vote Manipulation
2.2. Large Language Models
@@
Large Language Models (LLMs) have demonstrated significant advancements across a variety of tasks, including code generation (Zheng et al., 2023), static analysis (Li et al., 2023), and program repair (Wei et al., 2023). Currently, a set of prominent LLMs have been developed, such as the Generative Pre-trained Transformer (GPT) by OpenAI, LLaMA (Touvron et al., 2023) by Meta, and Gemini (Anil et al., 2023) by Google. These models possess extensive world knowledge, strong problem-solving capabilities, sophisticated reasoning skills, and a robust capacity for instruction following.
@@
LLMs exhibit significant emergent abilities (Wei et al., 2022a) that set them apart from traditional pre-trained models. They have three notable emergent abilities: in-context learning (Brown et al., 2020), instruction following (Ouyang et al., 2022), and step-by-step reasoning (Prystawski et al., 2023). For instance, small language models often struggle with complex tasks that require multi-step reasoning, such as analyzing the business logic concealed within smart contracts. In contrast, LLMs can adeptly manage these challenges using their emergent abilities. They utilize techniques like Chain-of-Thought (CoT) (Wei et al., 2022b) prompting to improve inference, thereby enabling them to address complex challenges effectively. In this paper, we utilize GPT-4-turbo as the foundational model in PromFuzz to analyze the hidden business logic in smart contracts. We focus on designing novel prompt strategies to trigger its emergent capabilities and implement practical measures to enhance its output robustness.
2.3. Motivating Example
We provide a functional bug discovered on Code4rena (Cod, 2024) as a motivating example, as shown in Figure 1. In the SimplePool contract, the transferFrom function facilitates the transfer of a specified amount of shares from the sender to the recipient using the balanceToShares function, which converts a specified token amount into shares based on the pricePerShare. However, a critical issue arises post-transfer: the approve operation erroneously subtracts the amountInShare instead of the amount. This mistake allows the approved spender potentially to transfer more tokens than intended, with the approved amount increasing in correlation with the pricePerShare.
A series of works, such as GPTScan (Sun et al., 2024) and SMARTINV (Wang et al., 2024), have explored the use of LLMs to detect functional bugs in smart contracts. GPTScan employs GPT-based matching analysis to initially identify potential functional bugs, which are then confirmed through static analysis. SMARTINV utilizes a fine-tuned LLaMa-7B model to infer invariants relevant to functional bug analysis and confirm bugs using validation algorithms. However, they are not consistently effective in identifying such bugs. Specifically, GPTScan cannot detect this issue as it fails during the GPT-based matching phase. SMARTINV generates invariants such as pricePerShare >  0 and balanceToShares(amount) >  0, which fail to capture the necessary details for analyzing this bug effectively.
PromFuzz introduces a novel approach that could effectively tackle and detect such bugs. Initially, PromFuzz employs an LLM-driven multi-perspective analysis that uses iterative questioning to pinpoint potential occurrences of this bug. Following this identification, PromFuzz extracts critical variables such as sender, allowances, and amount. An invariant checker is then constructed to verify the legitimacy of token quantity changes before and after transfers, which is integrated into the source code. Finally, PromFuzz involves bug-oriented fuzzing to ensure a comprehensive identification of functional bugs, completing the detection process.
@@ 2.3. Motivating Example
10 _approve(sender, _msgSender(), _allowances[sender][_msgSender()].sub(amountInShares,
11 "ERC20: transfer amount exceeds allowance"));
@@
13 }
14}
Figure 1. The Approval Not Clear bug.
3. PromFuzz Design
@@
Figure 2. Framework of PromFuzz.
3.1. LLM-driven Multi-Perspective Analysis
@@
Previous research has explored the use of LLMs to identify functional bugs in smart contracts by harnessing their significant emergent abilities. However, these studies often did not maximize the potential of LLMs, typically either inputting only high-level vulnerability descriptions or not fully engaging their capabilities. For example, Wang et al. (Wang et al., 2024) employed the LLaMA-7B model, but its relatively modest parameter size restricted its full emergent potential. Activating the emergent abilities of such models demands not only high-quality pre-training but also precise and effective instruction-following, which presents significant challenges. Furthermore, several previous works (Chen et al., 2023) (David et al., 2023) (Wang et al., 2024) have concentrated on enhancing the instruction-following capabilities of LLMs, frequently neglecting the crucial skill of step-by-step reasoning. This ability is vital for addressing complex problems, e.g., analyzing the obscured business logic of smart contracts, that require navigating through multiple layers of reasoning.
@@
Figure 3. Prompt template for the Auditor Agent.
@@
To address the above challenges, we design the Auditor Agent based on the CoT prompting strategy and the scenario and property matching methodology proposed by GPTScan (Sun et al., 2024). A “scenario” describes the code functionality under which a functional bug might occur, while a “property” details the vulnerable code attributes or operations. GPTScan initially selects ten representative functional bugs in smart contracts, following the classifications by Zhang et al. (Zhang et al., 2023). Upon further analysis, we observe that these categories could be streamlined for greater clarity and applicability.
Our refined approach consolidates the initial ten categories into four primary categories of functional bugs. Notably, we merge the categories of price manipulation by buying tokens and slippage into a single, more comprehensive category named non-AMM price oracle manipulation. Moreover, our manual analysis of the functional bugs leads to the identification of two new subcategories: improper handling of the deposit fee and wrong implementation of amount lock. We present the new classification results in Table 1.
Figure 3 outlines the prompt template for the Auditor Agent. First, GPT is informed that it is participating in a code auditing exercise. We then allow GPT to internally simulate the response formulation process five times and provide the most common answer to mitigate the randomness of the responses. Next, we input the target smart contract code and pose questions related to the functionality under which a functional bug might occur—i.e., the scenario. This question aids in identifying potential bug categories within the code. Finally, for any scenario where GPT confirms a potential issue, we follow up with specific questions about the vulnerable code attributes or operations—i.e., the property. If GPT responds affirmatively, we consider the presence of a functional bug; if not, we rule out that specific bug. The output of the Auditor Agent is fine-grained subcategories of bugs. Designing prompts requires only a one-time manual effort and can be easily extended to cover most functional bugs.
Figure 4. Prompt template for the Attacker Agent.
@@
To address the above challenges, we gather an extensive dataset of real-world attack incidents for each bug category. Through detailed manual analysis of these incidents, we extract critical information, e.g., distinctive bug features, to craft effective prompts based on the CoT prompt strategy. Bug features refer to attributes or behaviors in the code that present security risks under certain conditions, which could potentially be exploited under specific conditions to carry out malicious activities. The process of utilizing the Attacker Agent involves several key steps, outlined in Figure 4. First, similar to the Auditor Agent, GPT is informed that it is participating in an ethical hacking exercise. Next, the target smart contract code is inputted, and GPT is questioned about potential bug features that could be exploited for malicious activities. Finally, for bugs whose features have been recognized during the last question, we provide the corresponding attack model for these bugs as prompts to GPT. We then query GPT whether the identified features could lead to exploitable attacks based on the given attack model. An affirmative response from GPT indicates the presence of a functional bug, while a negative response helps us dismiss that particular bug. The output of the Attacker Agent is coarse-grained primary categories of bugs.
Figure 5. Auditor-Attacker results fusion process corresponds to the motivating example in Figure 1.
@@
Auditor-Attacker Results Fusion. To derive more accurate classifications of bugs, we integrate the results from both the Auditor Agent and the Attacker Agent. This fusion approach leverages the coarse-grained primary category insights from the Attacker Agent to refine and enhance the fine-grained subcategory determinations made by the Auditor Agent. The rules for results fusion are delineated as follows:
As shown in Algorithm 1, consider a scenario where for each analyzed code segment, the Attacker Agent identifies the primary bug categories $T$ at line 1, and the Auditor Agent identifies the subcategories $t$ at line 2. The fusion process operates under the following rules:
  1. 1)
    If $t$ is a subcategory of $T$, confirm $t$ as a candidate type (line 10-11).
  2. 2)
    In addition to confirming $t$, if $t$ belongs to $T$, also include all other subcategories of $T$ as candidate types to address potential overlaps or ambiguities (line 10-14).
  3. 3)
    If $t$ does not correspond to any subcategory of $T$, or if $T$ has not been identified, then exclude $t$ from the candidate types (line 17).
  4. 4)
    If $t$ is not determined but $T$ is identified, postpone the classification of this segment until further information is available or additional analysis is conducted (line 3).
For primary categories characterized by distinctly different subcategory features, such as insecure calculating logic and incorrect control mechanism, we apply rules 1), 3), and 4). For the other primary categories, we implement rules 2), 3), and 4). This selective application of fusion rules helps minimize false positives while maximizing the conversion of potential false negatives into true positives by the Auditor Agent. Although this method may initially increase the number of false positives, subsequent bug-oriented fuzzing processes are designed to effectively eliminate these inaccuracies.
Figure 5 illustrates an example of how our fusion rules are applied. First, the Auditor Agent identifies the code scenario as a token transfer and flags the property of this transfer as unsafe, classifying it under the subcategory of unauthorized transfer. Subsequently, the Attacker Agent analyzes potential bug features, indicating that these bugs could be exploited and categorizing them under the primary category of unauthorized behavior, which includes two subcategories: unauthorized transfer and approval not clear. Finally, we apply rule 2), 3) and 4) to integrate their analyses, leading to the identification of unauthorized transfer and approval not clear. The true positive identified in this code is approval not clear. Without this integration, we would fail to recognize this bug.
1 0:   $C$: smart contract code 0:   $B$: potential functional bug subcategory 1:  $T=Attacker\_Agent(C)$ 2:  $t=Auditor\_Agent(C)$ 3:  while $t\neq\varnothing$ do 4:     $t_{i}\leftarrow t.top()$ 5:     $t\leftarrow t-t_{i}$ 6:     while $T\neq\varnothing$ do 7:        $T_{j}\leftarrow T.top()$ 8:        $T\leftarrow T-T_{j}$ 9:        if $t_{i}$ is a subcategory of $T_{j}$ then 10:           $r_{mer}=Merge(T_{j},t_{i})$ 11:           $B=B.append(r_{mer})$ 12:           if  $j\in\{1,2\}$ then 13:              $r_{mut}=Mutation(T_{j},t_{i})$ 14:              $B=B.append(r_{mut})$ 15:           end if 16:        else 17:           Continue 18:        end if 19:     end while 20:  end while 21:  return  $B$
Algorithm 1 Auditor-Attacker Results Fusion
3.2. Invariant Checker Generation

### v2

# PromFuzz: Leveraging LLM-Driven and Bug-Oriented Composite Analysis for Detecting Functional Bugs in Smart Contracts

@@ Abstract
To address these challenges, we design and implement PromFuzz, an automated and scalable system to detect functional bugs in smart contracts. In PromFuzz, we first propose a novel Large Language Model (LLM)-driven analysis framework, which leverages a dual-agent prompt engineering strategy to pinpoint potentially vulnerable functions for further scrutiny. We then implement a dual-stage coupling approach, which focuses on generating invariant checkers that leverage logic information extracted from potentially vulnerable functions. Finally, we design a bug-oriented fuzzing engine, which maps the logical information from the high-level business model to the low-level smart contract implementations, and performs the bug-oriented fuzzing on targeted functions. We evaluate PromFuzz from $4$ perspectives on $5$ ground-truth datasets and compare it with multiple state-of-the-art methods. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs, marking at least a $50\%$ improvement in both metrics over state-of-the-art methods. Moreover, we perform an in-depth analysis on $10$ real-world DeFi projects and detect $30$ zero-day bugs. Our further case studies, the risky first deposit bug and the AMM price oracle manipulation bug on real-world DeFi projects, demonstrate the serious risks of the exploitable functional bugs in smart contracts. Up to now, $24$ zero-day bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling $\mathdollar 18.2$ billion from potential monetary losses.
I Introduction
@@
Blockchains have increasingly become a crucial component of the contemporary economic landscape. As of the writing of this paper, the total market capitalization of global cryptocurrencies has reached $\mathdollar 3.5$ trillion [3]. Smart contracts, which are fundamental pillars of the blockchain, facilitate the development of decentralized financial (DeFi) and various business transactions. Nevertheless, smart contracts are prone to exploitable bugs that often result in substantial financial losses. According to a report by CertiK, a leading Web3 security firm, over $\mathdollar 1.8$ billion was lost due to $751$ security incidents in 2023 [1].
@@
Currently, there are many works have focused on detecting exploitable bugs in smart contracts, which can be categorized into four main types based on their methodologies: static analysis [13, 38], fuzzing [10, 30], verification [32, 34], and symbolic execution [5, 31]. Despite these efforts, a recent study [45] reveals that over 80% of exploitable bugs remain undetected by current tools, which are primarily functional bugs. This limitation mainly stems from the tools’ reliance on simple or hard-coded oracles to identify exploitable bugs. Nevertheless, detecting functional bugs involves comprehending the high-level business logic before scrutinizing the low-level implementations in smart contracts. Therefore, there is a pressing need for a practical system that can automatically detect functional bugs in smart contracts.
I-A Challenges
@@ I-C Contributions
$\bullet$ To the best of our knowledge, we build the first ground-truth dataset for verified functional bugs in smart contracts, including 261 contracts from various DeFi projects. To facilitate future blockchain security research, we open-source both PromFuzz and the dataset at https://github.com/promfuzz.
$\bullet$ We have implemented and evaluated PromFuzz across $4$ dimensions using $5$ ground-truth datasets. The results show that PromFuzz achieves $86.96\%$ recall and $93.02\%$ F1-score in detecting functional bugs. PromFuzz demonstrates its superiority in at least a $50\%$ improvement in both metrics over state-of-the-art methods.
$\bullet$ Our extensive analysis of $10$ real-world DeFi projects demonstrates PromFuzz’s great performance in detecting real-world functional bugs. Up to now, PromFuzz has identified $30$ zero-day bugs in $6$ of the $10$ real-world DeFi projects, with $24$ of these bugs have been assigned CVE IDs. Our discoveries have safeguarded assets totaling $\mathdollar 18.2$ billion from potential monetary losses.
II Background
@@
II-A Functional Bugs in Smart Contract
In this paper, we primarily focus on four major categories of functional bugs, including a total of ten subcategories of functional bugs. These classifications cover the majority of functional bugs and are derived from prior research [33, 45].
@@ II-A Functional Bugs in Smart Contract
II-B Large Language Models
@@
Large Language Models (LLMs) have demonstrated significant advancements across a variety of tasks, including code generation [46], static analysis [20], and program repair [42]. Currently, a set of prominent LLMs have been developed, such as the Generative Pre-trained Transformer (GPT) by OpenAI, LLaMA [35] by Meta, and Gemini [7] by Google. These models possess extensive world knowledge, strong problem-solving capabilities, sophisticated reasoning skills, and a robust capacity for instruction following.
@@
LLMs exhibit significant emergent abilities [40] that set them apart from traditional pre-trained models. They have three notable emergent abilities: in-context learning [9], instruction following [27], and step-by-step reasoning [29]. For instance, small language models often struggle with complex tasks that require multi-step reasoning, such as analyzing the business logic concealed within smart contracts. In contrast, LLMs can adeptly manage these challenges using their emergent abilities. They utilize techniques like Chain-of-Thought (CoT) [41] prompting to improve inference, thereby enabling them to address complex challenges effectively. In this paper, we utilize GPT-4-turbo as the foundational model in PromFuzz to analyze the hidden business logic in smart contracts. We focus on designing novel prompt strategies to trigger its emergent capabilities and implement practical measures to enhance its output robustness.
II-C Motivating Example
We provide a functional bug discovered on Code4rena [2] as a motivating example, as shown in Figure 1. In the SimplePool contract, the function transferFrom is designed to transfer to the user the share tokens purchased with base tokens, while simultaneously deducting the corresponding amount from the user’s base token allowance. The auxiliary function balanceToShares converts a specified quantity of base tokens into share tokens according to the parameter pricePerShare, which defines the exchange rate between base tokens and share tokens. However, a critical flaw emerges from an incorrect subtraction operation at line 9 that allows users to acquire share tokens at an unfairly low cost. For instance, when pricePerShare = $1e18$, a user providing $5e18$ base tokens to the function balanceToShares at line 7 will receive 5 share tokens. These share tokens are then transferred through the function _transfer at line 8. While the conversion and transfer steps are correctly executed, the functional bug manifests in the subsequent deduction of payment. At line 9, the implementation erroneously deducts only 5 base tokens because the subtraction is applied to the variable amountInShare (equal to 5), rather than the correct variable amount (equal to $5e18$). The correct implementation should be updated as follows: _approve(sender,_msgSender(),_allowances[sender][_msgSender()].sub(amount,”ERC20: transfer amount exceeds allowance”));. As a result, the user obtains 5 share tokens for only 5 base tokens, a price significantly below the intended market rate. By repeatedly invoking transferFrom, a user can continue to exploit this functional bug to acquire additional share tokens at an unfairly low cost.
@@ II-C Motivating Example
9 _approve(sender, _msgSender(), _allowances[sender][_msgSender()].sub(amountInShares, "ERC20: transfer amount exceeds allowance"));
@@
Figure 1: The Approval Not Clear bug (line 9).
A series of works, such as GPTScan [33] and SMARTINV [37], have explored the use of LLMs to detect functional bugs in smart contracts. GPTScan employs GPT-based matching analysis to initially identify potential functional bugs, which are then confirmed through static analysis. SMARTINV utilizes a fine-tuned LLaMa-7B model to infer invariants relevant to functional bug analysis and confirm bugs using validation algorithms. However, they are not consistently effective in identifying such bugs. Specifically, GPTScan cannot detect this issue as it fails during the GPT-based matching phase. SMARTINV generates invariants such as pricePerShare >  0 and balanceToShares(amount) >  0, which fail to capture the necessary details for analyzing this bug effectively.
PromFuzz introduces a novel approach that could effectively tackle and detect such bugs. Initially, PromFuzz employs an LLM-driven multi-perspective analysis that uses iterative questioning to pinpoint potential occurrences of this bug. Following this identification, PromFuzz extracts critical variables, including sender, _allowances, and amount, to construct the invariant checker. The core idea for detecting the functional bug in the motivating example is to verify whether the user’s base token allowance before the payment, minus the required payment amount, equals the allowance after the payment. Specifically, the invariant checker records the user’s base token allowance before line 9 as old_allowance and inserts the invariant assertion, assert(old_allowance - amount == _allowances[sender][_msgSender()]);, immediately following line 9. Finally, PromFuzz involves bug-oriented fuzzing to ensure a comprehensive identification of functional bugs, completing the detection process.
III PromFuzz Design
@@
Figure 2: Framework of PromFuzz.
III-A LLM-driven Multi-Perspective Analysis
@@
In our dual-agent framework, the Auditor Agent and the Attacker Agent adopt distinct analysis strategies that differ in granularity. The intuition behind our design is that the two agents serve fundamentally different roles in the analysis pipeline. The Auditor Agent performs fine-grained inspections, dissecting the code at a detailed level to examine its semantics, business scenarios, and logical correctness. This mirrors real-world auditing practices, in which security professionals meticulously classify and validate individual program components to ensure reliability and security. Based on this fine-grained analysis, the Auditor Agent produces subcategories of bugs, offering a nuanced characterization of potential functional bugs. In contrast, the Attacker Agent conducts a coarse-grained analysis, focusing less on the accuracy of code understanding and more on the broader issue of exploitability. Rather than scrutinizing every line of code, the Attacker Agent seeks to determine whether certain functions, state transitions, or contract behaviors can be exploited to achieve unintended outcomes. Based on this coarse-grained analysis, the Attacker Agent outputs the primary categories of bugs, emphasizing the overarching attack surfaces that are most relevant for exploitation. This difference in analytical granularity reflects the agents’ fundamentally different perspectives on functional bug assessment: the Auditor Agent prioritizes accuracy and comprehensive understanding, while the Attacker Agent prioritizes feasibility of exploitation and its potential impact.
@@
Figure 3: Prompt template for the Auditor Agent.
@@
To address the above challenges, we design the Auditor Agent based on the CoT prompting strategy and the scenario and property matching methodology proposed by GPTScan [33]. A “scenario" describes the code functionality under which a functional bug might occur, while a “property" details the vulnerable code attributes or operations. GPTScan initially selects ten representative functional bugs in smart contracts, following the classifications by Zhang et al. [45]. Upon further analysis, we observe that these categories could be streamlined for greater clarity and applicability.
Our refined approach consolidates the initial ten categories into four primary categories of functional bugs. Notably, we merge the categories of price manipulation by buying tokens and slippage into a single, more comprehensive category named non-AMM price oracle manipulation. Moreover, our manual analysis of the functional bugs leads to the identification of two new subcategories: improper handling of the deposit fee and wrong implementation of amount lock.
Figure 3 outlines the prompt template for the Auditor Agent. First, GPT is informed that it is participating in a code auditing exercise. We then allow GPT to internally simulate the response formulation process five times and provide the most common answer to mitigate the randomness of the responses. Next, we input the target smart contract code and pose questions related to the functionality under which a functional bug might occur—i.e., the scenario. This question aids in identifying potential bug categories within the code. Finally, for any scenario where GPT confirms a potential issue, we follow up with specific questions about the vulnerable code attributes or operations—i.e., the property. If GPT responds affirmatively, we consider the presence of a functional bug; if not, we rule out that specific bug. Designing prompts requires only a one-time manual effort and can be easily extended to cover most functional bugs.
Figure 4: Prompt template for the Attacker Agent.
@@
To address the above challenges, we gather an extensive dataset of real-world attack incidents for each bug category. Through detailed manual analysis of these incidents, we extract critical information, e.g., distinctive bug features, to craft effective prompts based on the CoT prompt strategy. Bug features refer to attributes or behaviors in the code that present security risks under certain conditions, which could potentially be exploited under specific conditions to carry out malicious activities. The process of utilizing the Attacker Agent involves several key steps, outlined in Figure 4. First, similar to the Auditor Agent, GPT is informed that it is participating in an ethical hacking exercise. Next, the target smart contract code is inputted, and GPT is questioned about potential bug features that could be exploited for malicious activities. Finally, for bugs whose features have been recognized during the last question, we provide the corresponding attack model for these bugs as prompts to GPT. We then query GPT whether the identified features could lead to exploitable attacks based on the given attack model. An affirmative response from GPT indicates the presence of a functional bug, while a negative response helps us dismiss that particular bug.
@@
Auditor-Attacker Results Fusion. To derive more accurate classifications of bugs, we integrate the results from both the Auditor Agent and the Attacker Agent. This fusion approach leverages the coarse-grained primary category insights from the Attacker Agent to refine and enhance the fine-grained subcategory determinations made by the Auditor Agent. The Auditor-Attacker results fusion algorithm is provided in our supplementary material (§I).
III-B Invariant Checker Generation